OpenStudy (anonymous):

You have challenged four other students to come up with expressions equal to x3. Their submissions are below.

OpenStudy (anonymous):

where below ?

OpenStudy (anonymous):

Working on it hold tight please - thanks!!!

OpenStudy (anonymous):

Blake

OpenStudy (anonymous):

do you know how to multiply these things?

OpenStudy (anonymous):

@RED123url no i dont I can't find it in my lesson. I am struggling with division... there are four pics/student answers total so sit tight thanks

OpenStudy (ranga):

\[\Large x ^{m} \times x ^{n} = x ^{m+n}\]

OpenStudy (anonymous):

Zoe

OpenStudy (ranga):

\[\Large \frac{ x ^{m} }{ x ^{n}} = x ^{m-n}\]

OpenStudy (anonymous):

Eloise

OpenStudy (anonymous):

Dylan

OpenStudy (anonymous):

Okay thank you for your patience. Now, can I please get some kind of explaination on how to solve these problems? I'm so stuck. I'm not asking that you do my work for me, just help please. @ranga @RED123url

OpenStudy (anonymous):

I guess the guy did everything just follow the formulas he provided

OpenStudy (ranga):

I will do the first one: \[\Large x ^{\frac{ 4 }{ 3 }} \times x ^{\frac{ 7 }{3 }} = x ^{\frac{ 4 }{ 3 } + \frac{ 7 }{ 3 }} = x ^{\frac{ 11 }{ 3 }}\] \[\Large \frac{ x ^{\frac{ 11 }{ 3 }} }{ x ^{\frac{ 2 }{ 3}} } = x ^{\frac{ 11 }{ 3 } - \frac{ 2 }{ 3 }} = x ^{\frac{ 9 }{ 3 }} = x ^{3}\]

OpenStudy (anonymous):

What kind of a weird challenge is that? Anyway, also helpful to remember is that \(\sqrt[n]{x} = x^\frac{1}{n}\)

OpenStudy (anonymous):

@ranga Thank you. So, what you are showing in your response is that the first problem set up checks out and is correct to what the question is asking? And do you think you could just explain what the third and fourth ones are saying and how to go about them? That would REALLY help.

OpenStudy (anonymous):

@sacapuntas - thank you!!!

OpenStudy (ranga):

Yes the first one is properly set up and is correct. For the 3rd and 4th one use the formula that sacapuntas provided along with the two formulae I provided earlier. I will do the third one but this equation editor takes a long time.

OpenStudy (ranga):

\[\Large x ^{\frac{ 1 }{ 3 }}\times x ^{\frac{ 2 }{ 3}}\times x ^{\frac{ 4 }{ 3 }} \times \sqrt[3]{x} = x ^{\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }+\frac{ 4 }{ 3 }}\times x ^{\frac{ 1 }{ 3 }}\] \[\Large = x ^{\frac{ 7 }{3 }} \times x ^{\frac{ 1 }{ 3 }} = x ^{\frac{ 7+1 }{ 3 }} = x ^{\frac{ 8 }{ 3 }}\]

OpenStudy (ranga):

That was the fourth one not third.

OpenStudy (anonymous):

Thank you so much you are a life saver! @ranga

OpenStudy (ranga):

You are welcome.